# On gaps in the spectra of quasiperiodic Schrödinger operators with discontinuous monotone potentials

### v2

@@ Abstract.
We show that, for one-dimensional discrete Schrödinger operators, stability of Anderson localization under a class of rank one perturbations implies absence of intervals in spectra. The argument is based on well-known result of Gordon and del Rio–Makarov–Simon, combined with a way to consider perturbations whose ranges are not necessarily cyclic. The main application of the results is showing that a class of quasiperiodic operators with sawtooth-like potentials, for which such a version of stable localization is known, has Cantor spectra. We also obtain several results on gap filling under rank one perturbations for some general (not necessarily monotone) classes of quasiperiodic operators with discontinuous potentials.
@@ 1.1. Motivation and some of the main results
Let $f\colon\mathbb{R}\to[-\infty,+\infty)$ be a $1$-periodic function, and $H(x)$ be a family of discrete quasiperiodic Schrödinger operators on $\ell^{2}(\mathbb{Z})$:
@@
Two typical examples of $\gamma$-monotone potentials are $f(x)=\{x\}=x-\lfloor x\rfloor$ (the sawtooth potential) and $f(x)=\lambda\cot(\pi x)$, $\lambda>0$ (the Maryland model). We will say that $f$ is sawtooth-type if $f$ is $\gamma$-monotone and $f|_{[0,1)}$ extends to a (bounded) continuous map from $[0,1]$ to $\mathbb{R}$. Similarly, $f$ is Maryland-type if it is $\gamma$-monotone and $f|_{[0,1)}$ extends to a homeomorphism between $[0,1]$ and [,+][$\infty$,$\infty$][ - ∞ , + ∞ ].
@@
Let $f$ be sawtooth-type function and suppose that $\alpha$ is Diophantine. Then $\sigma(H(x))$ is a Cantor set ((((in particular, it contains no intervals)))).
@@
It is interesting to compare these results with those for Maryland-type potentials, which sometimes do not rely on localization and/or $\gamma$-monotonicity. We will call $f$ weak Maryland-type if $f|_{[0,1)}$ extends to a continuous map from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] and $f(0)=-\infty$, $f(1-0)=+\infty$. We will also call $f$ weak sawtooth-type if $f|_{[0,1)}$ extends to a (bounded) continuous map from $[0,1]$ to $\mathbb{R}$, with
$-\infty<f(0)<f(1-0)<+\infty.$
@@
The closure in the left hand side of (1.5) would not be necessary if one included $t=\infty$ into consideration by allowing infinite potentials. Again, we postpone the more general setting until Section 4 (see, in particular, Theorems 4.2 and 4.4), where this inclusion is performed systematically. In some situations, one can even prove that every point of $\mathbb{R}$ is an actual eigenvalue of some operator from our family, see Corollary 4.3. Note that, if one considers the union over all $t\in\mathbb{R}$ in (1.5), the result immediately follows from spectral averaging (for example, [17, Theorem 11.8]), which is why we assume $f(0)<f(1-0)$.
The general version of Theorem 1.2 also holds in higher dimensions for operators on $\mathbb{Z}^{d}$. A similar question about Theorem 1.1 appears to be more difficult. While the general argument behind Theorem 1.1 is abstract, it relies on two properties of the operator: Anderson localization that is stable under a family of rank one perturbations (that is, replacing $f$ by $f_{t}$) and the fact that a standard basis vector is cyclic on a sufficiently large subspace. In Theorem 1.1, both of these properties rely on the one-dimensional structure of the operators, the first one through [14], and the second one through Lemma 2.1 below. In an upcoming work, we intend to address the first question in a perturbative setting on $\mathbb{Z}^{d}$.
@@ 1.2. Stability of localization and absence of intervals
with real-valued potential $V$. Let also $P$ be a projection operator on the standard basis vector $e_{0}$: that is,
@@ 2. Absence of intervals in spectra
We will start from proving Theorem 1.4. Afterwards, we will establish its consequences for Schödinger operators with $\gamma$-monotone potentials, the most notable being Theorem 2.9 and Corollary 2.10.
@@
In the following lemma, $\mathbf{1}_{J}(H)$ denotes the spectral projector of a self-adjoint operator $H$ associated to an interval $J\subset\mathbb{R}$, and $e_{0}$ is the standard basis vector of $\ell^{2}(\mathbb{Z})$ associated to the origin. Suppose that $\varphi$ is an eigenfunction of $H$ with $\varphi(0)=0$. Then, the functions
$\varphi_{\pm}:=\varphi|_{\ell^{2}(\mathbb{Z}_{\pm})}$
@@
[…]
@@
Let $H$ be a discrete one-dimensional Schrödinger operator (1.6). Suppose that $J\subset\mathbb{R}$ is an open interval such that $\sigma(H)\cap J$ is pure point and $\mathbf{1}_{J}(H)e_{0}=0$ ((((in other words, every eigenfunction with eigenvalue in $J$ vanishes at the origin)))). Let $\varphi$, $\psi$ be two distinct normalized eigenfunctions with eigenvalues in $J$:
@@
As a consequence, the spectra of both operators in $J$ are purely point for $t\in[0,\varepsilon_{0})$. From Proposition 1.3, it follows that $\sigma(H_{c})$ cannot contain an interval. Let $I$ be a connected component of $J\setminus\sigma(H_{c})$. Then, all eigenvectors of $H$ with eigenvalues in $I$ must come from eigenvectors of $H_{0}$, which are orthogonal to $e_{0}$ and therefore vanish at the origin. From Corollary 2.2, $\sigma(H)$ cannot contain an interval in $I$, which completes the proof.  ∎
@@
Let $Q:=\mathbf{1}_{[\lambda_{1},\lambda_{2}]}(H)$ be the spectral projection of $H$ associated to the interval $[\lambda_{1},\lambda_{2}]$. Suppose that $Qe_{0}=0$, and that
@@
[…]
@@
Follows from applying Lemma 2.3 to intervals of the form $[\lambda+1/n,+\infty)\cap J$ and $(-\infty,\lambda-1/n]\cap J$. ∎
@@ 2.2. Quasiperiodic operators with $\gamma$-monotone potentials
where $\frac{p_{k}}{q_{k}}$ is the sequence of continued fraction approximants to $\alpha$. In addition to $\gamma$-monotonicity, we will also require
[…]
in order for the Lyapunov exponent of the operator family $H(x)$ to exist. The following results, in its final form, is established in [14].
@@
    $L(\cdot)$ is continuous in $E$ and one always has $L(E)\geq\log(\gamma/2)$. As a consequence, the results guarantee complete localization for large $\gamma$.
@@
    The integrated density of states is Lipschitz continuous. As a consequence, if the conclusion of the previous part holds, then for almost every $x$ the zero set of $L(E)$ does not contribute to spectral measure of $H(x)$. If $\beta(\alpha)=0$ (which is weaker than the Diophantine condition), then the previous claim implies localization for almost every $x$ for a class of $\gamma$-monotone potentials which includes sawtooth-type and Maryland-type.
@@
    Instead of $[0,1)$, one can also consider $\gamma$-monotone functions on $(0,1]$ (for example, by considering $-f(-x)$ instead of $f(x)$). As a consequence, each function from the family $f_{t}$ defined in (1.3) is $\gamma$-monotone in this sense.
@@
[…]
@@
Most of the claims of the lemma are known, especially if one avoids infinite coupling. We postpone the proof of the Lemma 2.8 until the end of Section 3, where these issues can be dealt with systematically. Note that the operator $H(x-0)$ (with the potential $V(n)=f(x+n\alpha-0)$) may not belong to the original family $\{H(x)\}_{x\in[0,1)}$. However, a natural way to extend the family is discussed in Section 4.
@@
Case 2: $y_{0}\in(0,1)$, $f(y_{0}-0)<f(y_{0})$. One can repeat the construction from Case 1 with $I=[f(y_{0}-0),f(y_{0})]$. Alternatively, one can consider $g(x)=-f(-x)$ as will be later done in Case 4.
@@
Case 4: $y_{0}=0$, $f(y_{0})=-\infty$, $f(1-0)<+\infty$. This case can be reduced to the previous case in two steps. First, replace $f(x)$ by $g(x):=-f(1-x)$. Denote the original operator by $H^{f}$, and the new operator by $H^{g}$. It is easy to see that $H^{f}(x)$ is unitarily equivalent to $-H^{g}(1-x)$, and $g$ is $\gamma$-monotone on $(0,1]$. Define the new function
@@
It is easy to see that $u(x)$ is now $\gamma$-monotone on $[0,1)$ and is within the class of operators considered in Case 2 (with $J$ replaced by $-J$). As a consequence, $\sigma(H^{u}(0))\cap(-J)$ cannot contain an interval. Since $H^{u}(0)$ is a finite rank perturbation of $H^{g}(1)$ and the latter is unitarily equivalent to $-H^{f}(0)$, we have that $\sigma(H^{f}(0))\cap J$ also cannot contain an interval. ∎
@@
As mentioned above, for non-generic $x$ the spectrum $\sigma(H(x))$ may contain isolated eigenvalues in the gaps of the Cantor set $\Sigma_{f}=\sigma_{\mathrm{ess}}(H(x))$. This would typically happen if, for some $x$, one has $f(x-0)\neq f(x)\neq f(x+0)$ (where one assumes +=$\infty$=$\infty$+ ∞ = - ∞). Theorem 4.2 provides a large class of such examples.
@@ 3. Operators with infinite potentials
Theorem 2.9 partially relies on Lemma 2.8. The proof of this lemma is fairly standard, but requires certain care when considering strong limits as the values of the potential approach infinity. A similar issue appears when one considers gap filling such as in Theorem 1.2: an operator with infinite coupling belongs to a family of rank one perturbations that are used to fill the gaps. In the proof of the gap filling result, we will need to use rational approximations and obtain several statements about continuity of spectra, based on the same ideas as in [3] and later developments, some of which were implemented in [16]. In the case of quasiperiodic operators with discontinuous/unbounded potentials, the usual issue is lack of compactness of the set of operators under consideration, either due to multiple ways of approaching a discontinuity in the limit, or due to the value of the potential approaching infinity.
In order to take full advantage of the above arguments, we will extend the class of operators under consideration, making the parameter space compact in the natural topology. This will include allowing the potential to take infinite values. For the convenience of the reader, we discuss the relevant theory (which we do not believe to be completely novel) in this section.
@@
  • The (generalized) spectrum of a Schrodinger operator $H$ is denoted by $\bar{\sigma}(H)$ and is a closed subset of $\overline{\mathbb{R}}$. We will always have $\sigma(H)=\bar{\sigma}(H)\cap\mathbb{R}$.

### v3

@@ Abstract.
We show that, for one-dimensional discrete Schrödinger operators, stability of Anderson localization under a class of rank one perturbations implies absence of intervals in spectra. The argument is based on well-known results of Gordon and del Rio–Makarov–Simon, combined with a way to consider perturbations whose ranges are not necessarily cyclic. The main application of the results is showing that a class of quasiperiodic operators with sawtooth-like potentials, for which such a version of stable localization is known, has Cantor spectra. We also obtain several results on gap filling under rank one perturbations for some general (not necessarily monotone) classes of quasiperiodic operators with discontinuous potentials.
@@ 1.1. Motivation and some of the main results
[…]
@@
Two typical examples of $\gamma$-monotone potentials are $f(x)=\{x\}=x-\lfloor x\rfloor$ (the sawtooth potential) and $f(x)=\lambda\cot(\pi x)$, $\lambda>0$ (the Maryland model). We will say that $f$ is sawtooth-type if $f$ is $\gamma$-monotone and $f|_{[0,1)}$ extends to a (bounded) continuous map from $[0,1]$ to $\mathbb{R}$. Similarly, $f$ is Maryland-type if it is $\gamma$-monotone and $f|_{[0,1)}$ extends to a homeomorphism between $[0,1]$ and $\infty$$\infty$].
@@
Let $f$ be sawtooth-type function and suppose that $\alpha$ is Diophantine. Then $\sigma(H(x))$ is a Cantor set ((in particular, it contains no intervals)).
@@
It is interesting to compare these results with those for Maryland-type potentials, which sometimes do not rely on localization and/or $\gamma$-monotonicity. We will call $f$ weak Maryland-type if $f|_{[0,1)}$ extends to a continuous map from $[0,1]$ to $\infty$$\infty$] and $f(0)=$-$\infty$, $f(1-0)=+\infty$. We will also call $f$ weak sawtooth-type if $f|_{[0,1)}$ extends to a (bounded) continuous map from $[0,1]$ to $\mathbb{R}$, with
$\infty<f(0)<f(1-0)<+\infty.$
@@
The closure in the left hand side of (1.5) would not be necessary if one included $t=\infty$ into consideration by allowing infinite potentials. Again, we postpone the more general setting until Section 4 (see, in particular, Theorems 4.2 and 4.4), where this inclusion is performed systematically. In some situations, one can even prove that every point of $\mathbb{R}$ is an actual eigenvalue of some operator from our family, see Corollary 4.3. Note that, if one considers the union over all $t\in\mathbb{R}$ in (1.5), the result immediately follows from spectral averaging (for example, [20, Theorem 11.8]), which is why we assume f(0)<f(10)f(0)<f(1-0).
The general version of Theorem 1.2 also holds in higher dimensions for operators on $\mathbb{Z}^{d}$. A similar question about Theorem 1.1 appears to be more difficult. While the general argument behind Theorem 1.1 is abstract, it relies on two properties of the operator: Anderson localization that is stable under a family of rank one perturbations (that is, replacing $f$ by $f_{t}$) and the fact that a standard basis vector is cyclic on a sufficiently large subspace. In Theorem 1.1, both of these properties rely on the one-dimensional structure of the operators, the first one through [16], and the second one through Lemma 2.1 below. The first question has been recently addressed in [5, 17].
@@ 1.2. Stability of localization and absence of intervals
with real-valued potential $V$. Let also $P$ be the orthohonal projection operator on the standard basis vector $e_{0}$: that is,
@@ 1.3. Structure of the paper
1.4. Acknowledgements
I. K. was supported by the NSF grants DMS–1846114, DMS–2052519, and the 2022 Sloan Research Fellowship. L. P. was supported by the EPSRC grant EP/V051636/1 and the Leverhulme Trust grant RPG-2023-325. R. S. was supported by the NSF grant DMS-2306327.
The authors would like to thank S. Jitomirskaya for valuable discussions and the anonymous referees for useful suggestions and remarks.
@@ 2. Absence of intervals in spectra
We will start from proving Theorem 1.4. Afterwards, we will establish its consequences for Schrödinger operators with $\gamma$-monotone potentials, the most notable being Theorem 2.10 and Corollary 2.11.
@@
In the following lemma, $\mathbf{1}_{J}(H)$ denotes the spectral projector of a self-adjoint operator $H$ associated to an interval $J\subset\mathbb{R}$, and $e_{0}$ is the standard basis vector of $\ell^{2}(\mathbb{Z})$ associated to the origin. Let
$\mathbb{Z}_{+}:=\mathbb{Z}\cap[0,+\infty),\quad\mathbb{Z}_{-}:=\mathbb{Z}\cap(-\infty,0].$
Suppose that $\varphi$ is an eigenfunction of $H$ with $\varphi(0)=0$. Then, the functions
$\varphi_{\pm}:=\varphi|_{\ell^{2}(\mathbb{Z}_{\pm})},$
@@
In our notation, $\|\cdot\|$ will always denote the $\ell^{2}$-norm.
@@
Let $H$ be a discrete one-dimensional Schrödinger operator (1.6). Suppose that $J\subset\mathbb{R}$ is an open interval such that $\sigma(H)\cap J$ is pure point and $\mathbf{1}_{J}(H)e_{0}=0$ ((in other words, every eigenfunction with eigenvalue in $J$ vanishes at the origin)). Let $\varphi$, $\psi$ be two distinct normalized eigenfunctions with eigenvalues in $J$:
@@
As a consequence, the spectra of both operators in $J$ are purely point for $t\in[0,t_{0})$. From Proposition 1.3, it follows that $\sigma(H_{c})$ cannot contain an interval. Let $I$ be a connected component of $J\setminus\sigma(H_{c})$. Then, all eigenvectors of $H$ with eigenvalues in $I$ must come from eigenvectors of $H_{0}$, which are orthogonal to $e_{0}$ and therefore vanish at the origin. From Corollary 2.2, $\sigma(H)$ cannot contain an interval in $I$, which completes the proof.  ∎
@@
Let $Q:=\mathbf{1}_{[\lambda_{1},\lambda_{2}]}(H)$ be the spectral projection of $H$ associated to the interval $\lambda_{1},\lambda_{2}]$. Suppose that $Qe_{0}=0$, and that
@@
Let $H$ be a discrete one-dimensional Schrödinger operator (1.6). Suppose that $J\subset\mathbb{R}$ is an open interval such that $\mathbf{1}_{J}(H)e_{0}=0$. Let $\varphi$ be an eigenfunction of $H$ with the eigenvalue $\lambda\in J$, and let $\theta\in\mathrm{Ran}\,\mathbf{1}_{J\setminus\{\lambda\}}(H)$. Then $\langle\theta_{+},\varphi_{+}\rangle=0$.
@@
Follows from applying Lemma 2.3 to intervals of the form $\lambda+1/n,+\infty)\cap J$ and $(-\infty,\lambda-1/n]\cap J$. ∎
@@ 2.2. Quasiperiodic operators with $\gamma$-monotone potentials
where $\frac{p_{k}}{q_{k}}$ is the sequence of continued fraction approximants to $\alpha$.
In the setting of one-dimensional Schrödinger operators, many results are stated in terms of the Lyapunov exponent $L(E)$, which is a non-negative real-valued function of energy associated to a one-dimensional ergdodic operator family. We will not be using the specific definition of $L(E)$ and would like to refer the reader to [15, 19, 16] in the setting of monotone potentials and to [6, Chapter 9,10] for a more general textbook treatment; the relevant properties are also summarized in Remark 2.7 below. In addition to $\gamma$-monotonicity, the condition
(2.6) $\int_{0}^{1}\log(1+|f(x)|)\,dx<+\infty,$
will be required in order for the Lyapunov exponent $L(E)$ of the operator family $H(x)$ to exist.
The following results, in its final form, is established in [16].
@@
    $L(\cdot)$ is continuous in $E$ and one always has $L(E)\geq\log(\gamma/2)$, see [15]. As a consequence, the conclusion of Proposition 2.6 guarantees complete localization for large $\gamma$.
@@
    The integrated density of states is Lipschitz continuous. As a consequence, if the conclusion of the previous part holds, then for almost every $x$ the zero set of $L(E)$ does not contribute to spectral measure of $H(x)$ (see [15, 19, 16]). If $\beta(\alpha)=0$ (which is weaker than the Diophantine condition), then the previous claim implies localization for almost every $x$ for a class of $\gamma$-monotone potentials which includes sawtooth-type and Maryland-type.
@@
    Instead of $[0,1)$, one can also consider $\gamma$-monotone functions on $(0,1]$ (for example, by considering f(x)-f(-x) instead of $f(x)$). As a consequence, each function from the family $f_{t}$ defined in (1.3) is $\gamma$-monotone in this sense.
@@
In order to state the second main result, we will need to be somewhat careful about the dependence of $\sigma(H(x))$ on $x$ as a set. Let $f$ be $\gamma$-monotone. We will call $x_{0}\in[0,1)$ generic if, for every $y_{0}\in(x_{0}+\alpha\mathbb{Z})\,\,\mathrm{mod}\,\,1$, we have one of the following:
@@
Most of the claims of the lemma are known, especially if one avoids infinite coupling. We postpone the proof of the Lemma 2.8 until the end of Section 3, where these issues can be dealt with systematically. Note that the operator H(x0)H(x-0) (with the potential $V(n)=f(x+n\alpha-0)$) may not belong to the original family $\{H(x)\}_{x\in[0,1)}$. However, a natural way to extend the family is discussed in Section 4.
Remark 2.9.
The claims of Lemma 2.8 also hold for operators defined on $\mathbb{Z}^{d}$, see Corollary 3.8 and the proof of Lemma 2.8 afterwards in Section 3.
@@
Case 2: $y_{0}\in(0,1)$, $f(y_{0}-0)<f(y_{0})$. One can repeat the construction from Case 1 with $I=[f(y_{0}-0),f(y_{0})]$. Alternatively, one can consider g(x)=f(x)g(x)=-f(-x) as will be later done in Case 4.
@@
Case 4: $y_{0}=0$, $f(y_{0})=$-$\infty$, $f(1-0)<+\infty$. This case can be reduced to the previous case in two steps. First, replace $f(x)$ by g(x):=f(1x)g(x):=-f(1-x). Denote the original operator by $H^{f}$, and the new operator by $H^{g}$. It is easy to see that $H^{f}(x)$ is unitarily equivalent to $H^{g}(1-x)$, and $g$ is $\gamma$-monotone on $(0,1]$. Define the new function
@@
It is easy to see that $u(x)$ is now $\gamma$-monotone on $[0,1)$ and is within the class of operators considered in Case 2 (with $J$ replaced by J-J). As a consequence, $\sigma(H^{u}(0))\cap(-J)$ cannot contain an interval. Since $H^{u}(0)$ is a finite rank perturbation of $H^{g}(1)$ and the latter is unitarily equivalent to $H^{f}(0)$, we have that $\sigma(H^{f}(0))\cap J$ also cannot contain an interval. ∎
@@
As mentioned above, for non-generic $x$ the spectrum $\sigma(H(x))$ may contain isolated eigenvalues in the gaps of the Cantor set $\Sigma_{f}=\sigma_{\mathrm{ess}}(H(x))$. This would typically happen if, for some $x$, one has $f(x-0)\neq f(x)\neq f(x+0)$ (where one assumes $\infty$$\infty$). Theorem 4.2 provides a large class of such examples.
@@ 3. Operators with infinite potentials
In this section, we work with operators (3.1) acting on $\ell^{2}(\mathbb{Z}^{d})$. We will prove Lemma 2.8, as well as lay some groundwork for proving Theorem 1.2 in higher dimensions. Note that neither of these results rely on localization and therefore do not involve $L(E)$ or $\beta(\alpha)$ that are only defined in dimension one anyway.
The proof of Lemma 2.8 is fairly standard, but requires certain care when considering strong limits as the values of the potential approach infinity. A similar issue appears when one considers gap filling such as in Theorem 1.2: an operator with infinite coupling belongs to a family of rank one perturbations that are used to fill the gaps. In the proof of the gap filling result, we will need to use rational approximations and obtain several statements about continuity of spectra, based on the same ideas as in [3] and later developments, some of which were implemented in [19]. In the case of quasiperiodic operators with discontinuous/unbounded potentials, the usual issue is lack of compactness of the set of operators under consideration, either due to multiple ways of approaching a discontinuity in the limit, or due to the value of the potential approaching infinity.
In order to take full advantage of the above arguments, we will extend the class of operators under consideration, making the parameter space compact in the natural topology. This will include allowing the potential to take infinite values. For the convenience of the reader, we discuss the relevant theory, which we do not believe to be completely novel. The results of this section are obtained either for general Schrödinger operators on $\ell^{2}(\mathbb{Z}^{d})$ or, as in Corollary 3.8, for general monotone quasiperiodic operators on $\ell^{2}(\mathbb{Z}^{d})$ with rationally independent frequency vectors.
@@
  • The (generalized) spectrum of a Schrödinger operator $H$ is denoted by $\bar{\sigma}(H)$ and is a closed subset of $\overline{\mathbb{R}}$. We will always have $\sigma(H)=\bar{\sigma}(H)\cap\mathbb{R}$.